OpenStudy (xaze):

Help me out, please? :3 Which is a factor of 15xy - 45x - 6y + 18?

OpenStudy (xaze):

It seems pretty easy but im not sure how to do it >n<

OpenStudy (anonymous):

15xy – 45x – 6y + 18 = 3 (5xy - 15x - 2y + 6) = 3 ( 5x(y - 3) - 2 (y - 3) ) = 3 (5x-2) (y - 3).

OpenStudy (mrnood):

look at all th terms separately can you see a number which divides into all of them? 15xy 45x 6y 18 (hint - look at the last one first)

OpenStudy (xaze):

Hmmm. 3?

OpenStudy (mrnood):

yes well done - that is 1 factor (your question seems to ask for just one factor)

OpenStudy (anonymous):

can i get fan

OpenStudy (xaze):

Sure, dude ^^

OpenStudy (anonymous):

i am not a dude

OpenStudy (anonymous):

=)

OpenStudy (xaze):

@MrNood, but the answers are shown as "5x-3" and "y-6" and stuff >.<

OpenStudy (xaze):

@joejoe213 Ah, my apologies, ma'am. ^^ Yes, I will fan you. xD

OpenStudy (anonymous):

its ok

OpenStudy (xaze):

So based on joejoe213's answer, 3 (5x-2) (y - 3), (5x-2) is one of the choices. So I believe (5x-2) is the correct answer.

OpenStudy (mrnood):

that is correct - but you didn't post the whole options. 3 is a factor as requested